OpenStudy (stormoshawty):

A recent Daytona 500 race was won by Dale Earnhardt, Jr. He completed the 500 mile race in 3.2 hours. What was his average racing speed?

OpenStudy (anonymous):

miles per hour per means divided by \[\frac{\text{ number of miles}}{\text{number of hours }}=\frac{5000}{3.2}\]

OpenStudy (anonymous):

oops!!\[\frac{500}{3.2}\]

OpenStudy (anonymous):

it is the daytona five hundred not the daytona five thousand

OpenStudy (anonymous):

got a calculator?

OpenStudy (stormoshawty):

Yes lmao

OpenStudy (anonymous):

ok then you should be in good shape i use this http://www.wolframalpha.com/input/?i=500%2F3.2